baseball outs. In a game, 12 of a baseball team's 27 outs were fly balls. 25 percent of the outs made by infielders and 100 percent of the outs made by outfielders were fly balls.
a. Write a linear systen=m you can use to find the number of outs made by infielders and the number of outs made by outfielders.
b. Graph your linear system
c. How many outs were made by infielders? How many outs were made by outfielders?

Answer by Theo(13342)

x = number of outs made by infielders.
y = number of outs made by outfielders.

total number of outs = 27, so x + y = 27.

25% of outs made by infielders and 100% of outs made by outfielders were fly balls, so .25x + y = 12

your two equations that need to be solved simultaneously are:

x + y = 27
.25x + y = 12

subtract the second equation from the first to get .75x = 15

divide both sides of the equation by .75 to get x = 20

number of outs made by infielders is 20.
number outs made by outfielders is 7.

.25 * 20 = 5 outs made by infielders that were fly balls.
1 * 7 = 7 outs made by outfielders that were fly balls.

total outs by infielders and outfielders is 27.
total fly outs made by infielders and outfielders is 12.

solution is confirmed to be good.

solution is 20 outs made by infielders and 7 outs made by outfielders.